OpenStudy (phagey):

A large bucket that is a full of water has a small leak on the bottom. The bucket loses water at the rate of 0.5 gallons per min. After 6 mins the bucket contains exactly 9 gallons of water. How many gallons of water were initially in the bucket? Also write an equation in point slope form to model the number of gallons (y) of water in the bucket after x minutes

OpenStudy (phagey):

@Michele_Laino

OpenStudy (michele_laino):

hint: if I call with \(y_0\) the initial amount of water of the bucket, and with \(x\) the number of minutes, then, after \(x\) minutes the amount of water, is: \(y_0-0.5 \cdot x\)

OpenStudy (michele_laino):

now, if we specialize at the case \(x=6\), then we can write: \[\huge {y_0} - 0.5 \cdot 6 = 9\] please solve for \(y_0\)

OpenStudy (phagey):

i got 12 @Michele_Laino

OpenStudy (phagey):

@Michele_Laino

OpenStudy (michele_laino):

ok! So, the initial amount of water is \(12\) gallons Now, after \(x\) minutes, the remaining amount of water, will be: \(12-0.5 \cdot x\), since we have to start from the initial amount which is \(12\) gallons

OpenStudy (michele_laino):

so, what is the equation in the point slope form ?

OpenStudy (phagey):

i really dont know @Michele_Laino

OpenStudy (michele_laino):

it is simple, the generic amount of water contained by the bucket after \(x\) minutes, is the dependent variable \(y\), so the requested equation, is: \(y=12-0.5x\)

OpenStudy (phagey):

ohhh i feel so stupid thank you @Michele_Laino

OpenStudy (michele_laino):

thanks!! :)